OpenStudy (anonymous):

someone help explain this to me please the Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x

OpenStudy (anonymous):

Function: f(x) = 4x+8 Inverse: g(x) = x+8 / 4

OpenStudy (anonymous):

how do i do this? please someone explain

OpenStudy (anonymous):

i dont need the answers im not trying to cheat i just need help

OpenStudy (loser66):

let y = 4x +8 , now, change the position of x and y, what do you get?

OpenStudy (anonymous):

x = 4y + 8

OpenStudy (loser66):

ok, now isolate your "new" y by subtract both sides by 8 and then divide both sides by 4, what do you get for that "new" y

OpenStudy (anonymous):

x = 4y + 8 -8 -8 =-4

OpenStudy (anonymous):

now i divide right

OpenStudy (loser66):

where is your y and x?